# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v1

@@ 3.2 Instruction Supervised Fine-Tuning
As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of 50 linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model’s output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix A.1.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some idioms, the final distribution is approximately 70:30 in favor of files with no violations. This results in a total of 53k training instances spanning 50 idioms.
@@ 3.3 Verifiable Reward Model and Preference Optimization
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter’s line numbers (yellow circle in “Verifiable Reward Model”, Figure 1) as ground truth and the model’s predicted lines (blue circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$.
Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
@@ 3.4 Training with Reasoning Traces
Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning. To obtain CoT traces that help the LLM arrive at the correct answer, we adopt a rejection sampling approach (“Rejection Sampling SFT” in Figure 2) for collecting SFT data. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained LLM capable of CoT reasoning (e.g., Qwen3-4B), and compute the reward $r_{y_{i}}$ for each response, following a similar procedure as in RS-DPO. However, instead of forming contrastive pairs, we discard any $y_{i}$ with reward $r_{y_{i}}<\gamma$, where $\gamma=1$. In other words, we discard CoT–response pairs that are incorrect or improperly formatted and retain only those that are fully correct. The reward is applied only to the final response, obtained after extracting or parsing the CoT trace. We also remove any cases where the CoT fails to terminate or produce a final answer.
For some inputs $x$, no valid $y_{i}$ may be found. In such cases, we skip the input entirely. To promote meta-task diversity, we retain at most two valid responses $y_{i}$ per input. We sample multiple $y_{i}$ only for inputs with at least one violation, and only a single $y_{i}$ for inputs with no violations, this helps maintain a similar no-violation-to-violation ratio of 71:29 as in the regular SFT data. The latter case is also more likely to yield a valid CoT–response pair within the token limit. When more valid responses are available than needed, we keep the shortest $y_{i}$ completions, as they typically reflect more concise reasoning (the final answer usually consumes a similar number of tokens across samples). Following this policy, we collect 52.7k instances, which we use to train the base model using SFT.
After training the LLM on this dataset, we obtain a CoT-capable SFT model $\Phi_{CoT}^{SFT}$. We then apply the same RS-DPO procedure described in Section 3.3, with the only modification being that each $y_{i}$ now contains both the CoT trace and the final response.
@@ 4.1 PEP Hard Idiom Benchmark
To evaluate whether MetaLint training helps the base LLM learn to interpret high-level idiom specifications, and apply this skill to harder, more abstract idioms that linters struggle to detect, we construct a benchmark of “hard idioms” sourced from PEPs. We begin by identifying fifteen PEPs that define semantic or abstract behaviors beyond what syntax or token patterns capture, then design specific heuristics for each to detect code that violates their guidelines. Using these heuristics, we search the STACK‑V2 corpus for violations, prioritizing high recall over strict precision to retrieve a broad set of candidates for manual selection. These “hard idioms” cannot be reliably detected by simple pattern matching, making them ideal for testing whether models truly understand idiom specifications or memorize frequent patterns seen in the Ruff idioms. From the retrieved examples, we handpick 15–20 representative files per PEP that clearly illustrate each violation. For each file, we identify the exact lines showing the violation (“before” snippet) and author a corrected version (“after” snippet) that complies with the PEP, forming the ground truth. To balance the benchmark, we include negative examples (code without violations) by selecting files retrieved for one PEP and confirming they don’t violate others. This results in 536 challenging examples, with 52% containing at least one violation and 48% being violation-free, allowing us to evaluate MetaLint’s ability to generalize from easy to hard idioms.
4.2 Evaluation Metrics
We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs. For both tasks, we compute precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each idiom, treating each idiom as a distinct class. Since the distribution of idiom occurrences may be imbalanced, we use macro averaging across idioms to weigh each class equally.
Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations. For each instance $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$, we define the indicator variable:
The idiom detection metrics per idiom $I$ are defined as:
We compute the final detection metrics by macro averaging across all idioms:
For idiom localization, we compute set level precision, recall, and F1-score for each instance $(x,y)\in\mathcal{D}$, comparing the predicted and ground truth violation lines:
4.3 Baselines
@@
We evaluate a range of state-of-the-art open and closed-source LLMs, trained for coding and reasoning tasks, on the PEP hard idiom benchmark to contextualize the improvements achieved by the MetaLint framework in flagging idiom violations from high-level specifications. Among open source models, we include instruction-tuned variants from the Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), and Qwen3 (Team, 2025) families. For closed-source models, we evaluate several OpenAI systems, including GPT-4o (Hurst et al., 2024), o3 mini, and o4 mini (OpenAI, 2025), and GPT-4.1 (OpenAI, 2025). We choose these models for their strong performance on a variety of code and reasoning benchmarks, and we evaluate whether code-specific training leads to measurable improvements. We also compare different model sizes, ranging from 3B and 4B to 32B parameters, in the open-source group
4.4 Evaluating Transfer Performance
To evaluate the extent to which the MetaLint training framework encourages adaptive reasoning over idiom specifications, as opposed to memorization, we construct a synthetic test set of 5.3k instances spanning 50 Ruff idioms, using the same data generation procedure described in Section 3.1. The data has a similar distribution of no-violation to violation examples as the SFT training set, with a 74:26 split in favor of no violations. We carefully select idioms to vary in their degree of overlap with those used for training, as shown in Figure 3. The selected idioms fall into three categories:
In-Domain: We select 5 idioms that are identical to those seen during SFT training, to assess whether MetaLint improves performance on idioms it was explicitly trained on.
Near Transfer: We include 10 idioms with specifications that closely resemble those in the training set but differ slightly. This category is crucial for probing memorization: if MetaLint relies too heavily on memorized training idioms, it may misclassify or underperform on these similar-looking test idioms due to interference.
Far Transfer: We choose 35 idioms that are substantially different from the training idioms to assess whether the LLM can rely solely on the provided specification. Success in this category would indicate that MetaLint fosters generalizable reasoning skills rather than dependence on specific training instances.
4.5 Evaluating Easy-to-Hard Generalization
To evaluate easy-to-hard generalization, we use the PEP hard idiom benchmark and compare the MetaLint-trained model against the baselines described in Section 4.3. The PEP idioms are held out for all models, including MetaLint, which is trained only on synthetic data generated from Ruff idioms. If training on the easier Ruff idioms leads to improved performance of MetaLint over the base models on the harder PEP idioms, it provides evidence that MetaLint encourages adaptive reasoning rather than memorization, thereby supporting easy-to-hard generalization.
@@ 5 Results
To test whether MetaLint enables the base LLM to interpret idiom specifications rather than simply memorize training examples, we evaluate its ability to generalize beyond the 50 Ruff idioms seen during training (Section 4.4) to both similarnear-transfer and significantly different far-transfer idioms. We also hypothesize that the RS-DPO (preference optimization) stage improves generalization over SFT alone, and that chain-of-thought (CoT) variants outperform their standard counterparts. Additionally, we assess easy-to-hard generalization by training on synthetic data for simpler Ruff idioms and evaluating on harder PEP idioms; if the model has truly learned to interpret high-level specifications, it should improve on these harder cases as well. Once again, we expect RS-DPO and CoT variants to yield stronger generalization. We present and analyze the results of these experiments in the following sections.
5.1 Ruff Idioms Performance
Table 1 presents Qwen3-4B’s idiom detection and localization performance on the full Ruff idiom transfer test set (Section 4.4) after MetaLint training, comparing CoT and non-CoT variants across both SFT and DPO stages. The untrained Qwen3-4B without CoT achieves decent detection precision but performs poorly on detection recall and localization. After SFT training, it shows substantial improvements in precision and all localization metrics, with modest recall gains. RS-DPO training further boosts detection recall and localization significantly (over 3× and nearly 5×, respectively), although with a slight drop in detection precision.
This precision drop stems from contrastive training examples where input $x$ has no violations, reducing the model’s accuracy in identifying such cases. We observe a strong correlation between the fraction of no-violation samples in training and detection precision. Including more such samples can improve precision but at the cost of recall and localization, leading to an unfavorable tradeoff. We find that including just 0.05% of no-violation data yields the best balance, minimizing precision loss while maximizing recall and improving detection F-score. Detailed ablation results are reported in Appendix A.2.2.
For the CoT variant, the untrained model already achieves strong detection precision and recall, along with high localization metrics—outperforming the SFT-only non-CoT model and trailing only slightly behind the SFT+RS-DPO non-CoT model (except in precision). To train the CoT model, we first apply RS-SFT to promote correct CoT reasoning and response formatting. This step is critical, as direct RS-DPO training alone causes metric collapse due to lack of absolute correctness supervision (since RS-DPO optimizes relative preference). RS-SFT leads to similar gains as non-CoT SFT, improving precision, recall, and localization. Final RS-SFT+RS-DPO training with 5% no-violation data mirrors the trends seen in the non-CoT case: increased recall, slight drop in precision, and overall gains in detection F-score and localization. Overall, the CoT model (RS-SFT+RS-DPO) is the strongest, achieving the highest precision and localization performance. However, the non-CoT SFT+RS-DPO model closely matches it in recall and approaches it on localization. Given that CoT inference is nearly 6× slower111measured on vLLM with 1× A100 80GB GPU the non-CoT model may be preferable in settings where detection recall and localization are the primary concerns.
5.2 Ruff Idioms Transfer Learning Performance
Table 2 presents MetaLint’s idiom detection performance across training setups for each transfer condition (in-domain, near-transfer and far-transfer) on the Ruff idiom benchmark. Untrained models perform poorly on in-domain idioms, especially in recall, suggesting these idioms are more subtle and difficult. Their stronger baseline performance on near and far transfer tasks likely reflects simpler or more surface-level idioms in those settings. SFT improves both precision and recall in-domain, but only precision improves in transfer settings, indicating limited generalization. Adding RS-DPO leads to consistent gains across all conditions, with particularly strong improvements in near and far transfer. This suggests that preference tuning encourages models to prioritize patterns that generalize beyond static rules.
CoT Untrained models already show high precision, likely due to the structured reasoning helping with correct interpretation of the idiom specifications and the output formatting instructions. RS-SFT improves recall with minimal changes to precision, though a small recall drop in near transfer may be due to a high baseline level of performance. Adding RS-DPO further boosts recall in all settings while maintaining strong precision. The largest recall improvements are seen in the in-domain setting, where idioms are hardest. These results offer some support for our claims that instruction-tuned models trained on synthetic linter signals generalize to novel idioms (the near and far transfer settings).
5.3 PEP Hard Idiom Benchmark Performance
Table 3 presents the results of evaluating the MetaLint DPO models and all baselines on the PEP hard idiom benchmark (a more comprehensive set of results is shown in Table 7). The MetaLint Qwen3-4B models significantly outperform the base Qwen3-4B models across all metrics in both the CoT and non-CoT settings, for both idiom violation detection and localization.
Interestingly, the non-CoT model achieves significantly higher detection recall and slightly higher F-score than the CoT variant, despite being less precise. Our analysis reveals that the CoT model’s lower detection recall results from its more conservative interpretation of the idiom specification and errors such as incorrect idiom interpretation, overthinking, and skipping lines, as detailed in Appendix A.3.3. In terms of detection F-score, the MetaLint models perform competitively, with only the closed-source GPT models, Qwen3-32B with CoT, and DeepSeek-R1-Distill-Qwen-32B with CoT surpassing the non-CoT model. Notably, the non-CoT model achieves the highest detection recall among all models evaluated, while the CoT model ranks among the best in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
For localization, the MetaLint models trail behind larger 32B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix A.3.2). This is notable given that the MetaLint models are much smaller (4B parameters) and trained only on synthetic data derived from easy idioms.
Overall, the strong performance, especially the best-in-class recall of the non-CoT model, demonstrates the effectiveness of our framework in achieving easy-to-hard generalization. This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction tuning, rather than relying on rote memorization.

### v2

@@ 3.2 Instruction Supervised Fine-Tuning
As discussed in Section 3.1, we train the target LLM $\Phi$ on a set of linter-detectable, easy idioms $I_{\mathcal{L}}$, using the corresponding meta-task specifications $M_{I_{\mathcal{L}}}$ and a set of source code files $\mathcal{F}$. The input to the model consists of a prompt $p$, which combines a meta-task specification $M_{I}$ for some $I\in I_{\mathcal{L}}$ with a source code file $f\in\mathcal{F}$. The model's output is a list of idiom violations in the file, denoted as $V_{f,I}$, formatted as a JSON list with one violation per line (see example output in Appendix C.1). In cases where there are no violations ($|V_{f,I}|=0$), the model is expected to output the phrase NO VIOLATIONS FOUND. We attempt to balance the data between positive (violations) and negative (no violations) examples as much as possible; however, due to the rarity of some Python idioms, the final distribution is approximately 70:30 in favor of files with no violations for Python data, but roughly 53:47 (PMD) and 50:50 (JEP Tree-Sitter) for the Java data. This results in a total of 53k synthetic training instances spanning 50 idioms (a subset of all the idioms detectable by Ruff) for Python Ruff data and 96.8k instances spanning 269 idioms for Java PMD and 127.3k instances spanning 15 idioms for tree-sitter data, respectively.
@@ 3.3 Verifiable Reward Model and Preference Optimization
Reward Model Design: The reward model evaluates model outputs by comparing predicted violations against those flagged by the linter, treating the linter's line numbers (blue circle in ``Verifiable Reward Model'', Figure 1) as ground truth and the model’s predicted lines (yellow circle) as predictions. Reward is computed using set-based precision, recall, and F1-score (visualized via the Venn diagram in the same figure), based on line-level overlap. Since each meta-task $M_{I}$ corresponds to a single idiom $I$, we compute one F1-score (reward) per instance.
Sampling Contrastive Pairs: We begin with an SFT policy model $\Phi^{SFT}$ and sample $k=5$ outputs $y_{i}$, $i\in\{1,\dots,k\}$ for each input $x$, using a range of temperature values $\tau=\{0,0.3,0.5,0.7,1.0\}$ to promote output diversity. Each response $y_{i}$ receives a reward $r_{y_{i}}$, and for each pair $(y_{i},y_{j})$, we compute the reward gap $|r_{y_{i}}-r_{y_{j}}|$. Pairs with a gap greater than the threshold $\eta=0.2$ are added to the preference dataset $\mathcal{D}_{p}$. For any such pair where $r_{y_{i}}\geq r_{y_{j}}+\eta$, we assign $y_{win}=y_{i}$, $y_{lose}=y_{j}$, and store the instance $(x,y_{win},y_{lose})\in\mathcal{D}_{p}$. Following Khaki et al. (2024), we train the preference-tuned model $\Phi^{RL}$ using the DPO objective:
@@ 3.4 Training with Reasoning Traces
Finally, inspired by the success of reasoning-augmented models in code and math tasks, and their demonstrated effectiveness in improving CWE detection performance in LLMs (Khare et al., 2023), we propose SFT and DPO methods that incorporate chain-of-thought (CoT) reasoning. To obtain CoT traces that guide the LLM to correct answers, we adopt a rejection sampling approach (``Rejection Sampling SFT'' in Figure 2) for SFT data collection. For each input $x$, we sample $k=5$ responses $y_{i}$, $i\in\{1,\dots,k\}$, from a base untrained CoT-capable LLM (e.g., Qwen3-4B), and compute a reward $r_{y_{i}}$ for each, following the RS-DPO procedure in Figure 2. Instead of forming contrastive pairs, we discard any $y_{i}$ with $r_{y_{i}}<\gamma$, where $\gamma=1$, i.e., CoT–response pairs that are incorrect or improperly formatted. Rewards are applied only to the final response, obtained after parsing the CoT trace, and we also remove cases where the CoT fails to terminate or yield an answer. If no valid $y_{i}$ is found for an input $x$, we skip it. To promote meta-task diversity, we retain at most two valid responses per input: multiple $y_{i}$ only for violation cases and a single $y_{i}$ otherwise. This maintains the 71:29 no-violation-to-violation ratio of Ruff Python SFT data, with the latter more likely to fit within token limits. When excess valid responses exist, we keep the shortest completions, as they typically reflect more concise reasoning (final answers are of similar token length across samples). Following this policy, we collect 52.7k Python training instances from Ruff data, which we use to train the reasoning-enabled base Qwen3-4B with SFT. This yields a CoT-capable SFT model $\Phi_{CoT}^{SFT}$ for Python code quality analysis. We then apply the RS-DPO procedure in Section 3.3 and Figure 2, with the only change being that each $y_{i}$ now includes both the CoT trace and final response.
Figure 1: MetaLint: (1) Synthetic data generation with linters/tools, (2) Supervised Instruction Fine-Tuning (SFT) on this data, and (3) Verifiable Reward Model derived from the linter.
Figure 2: MetaLint: Preference Optimization using reward model: (4) Rejection Sampling Direct Preference Optimization (RS-DPO), and (5) Rejection Sampling Supervised Fine-Tuning (RS-SFT).
@@ 4 Experiments
4.1 Evaluation Metrics
We evaluate the LLM’s ability to detect idiom violations through two tasks: detection, which assesses whether a given idiom is violated in a code file, and localization, which evaluates whether the model accurately identifies the specific line numbers where the violation occurs. For both tasks, we report precision, recall, and F-score metrics. Detection metrics are calculated at the corpus level for each idiom, treating each as a separate class, while localization metrics are computed at the instance level using set-based precision, recall, and F-score for the ground truth and predicted sets of violating line numbers. To handle potential class imbalance, we use macro-averaging across idioms and exclude NO VIOLATION as a class to penalize models that only predict NO VIOLATIONS FOUND (such models will score zero on all detection metrics). For localization, metrics are averaged only across instances with at least one line of idiom violation in the ground truth. Details of the formal definitions and exact computations of precision, recall, and F-scores for detection and localization are provided in Appendix D.1.
4.2 Generalization on Synthetic Data
To evaluate whether MetaLint training produces adaptive LLMs that handle evolving best practices and novel idioms at test time, we explore transfer settings spanning Python & Java.
Ruff Python Idioms: We construct a 5.3k-instance synthetic test set spanning 50 Ruff idioms, using the data generation procedure from Section 3.1. The data has a 74:26 no-violation-to-violation split, similar to the SFT training set. Idioms are chosen to vary in overlap with training idioms (Figure 3) and fall into three categories:
In domain. 5 idioms identical to those in SFT training, to assess whether MetaLint improves performance on explicitly trained idioms.
Near transfer. 10 idioms with specifications similar but not identical to training idioms to probe memorization. Reliance on memorized patterns, may hurt performance due to interference.
Far transfer. 35 idioms distinct from training, to test whether the LLM can follow the provided specification and adapt to novel idioms at test time.
For these experiments, we use Qwen3-4B (with and without reasoning) and Llama-3.2-3B-Instruct to study the effect of test-time compute and model family.
PMD and JEP Tree-Sitter Idioms: For Java, we construct two synthetic test sets: 5.1k instances (54:46 split) spanning 269 PMD idioms, and 6.4k instances (50:50 split) spanning 15 JEP idioms (Table 5), flagged via tree-sitter queries. We evaluate in-domain performance by training the base LLM on the corresponding training set (Section 3.2), and also study transfer between PMD and JEP idioms to test adaptation to novel Java idioms. These experiments use Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct to assess the effect of model scale.
@@ 4.3 PEP Hard Idiom Benchmark
Benchmark Construction: To test whether MetaLint helps LLMs interpret high-level idiom specifications and generalize to nuanced idioms that linters miss, we construct a benchmark of ``hard idioms'' from 15 PEPs defining semantic or abstract behaviors beyond syntax. We design heuristics per PEP (Table 13, 14 and 15) to detect guideline violations and search the STACK-V2 corpus, prioritizing recall to retrieve broad candidate sets for manual selection. These idioms cannot be reliably detected by simple pattern matching, making them ideal for evaluating model's true understanding versus rote memorization. From the candidates, we handpick 15–20 representative files per PEP, and annotate the precise line ranges (“start” and “end”) of the violated code, providing ground-truth for localization. We add negative examples for each PEP by picking files retrieved for a different PEP and making sure the current PEP is not violated, in order to have a balanced distribution of violation and no-violation cases. The final benchmark contains 536 examples (52% violations, 48% violation-free), enabling evaluation of MetaLint’s generalization from easy to hard Python idioms.
Evaluating Easy-to-Hard Generalization: We use the PEP hard idiom benchmark to test whether training on synthetic data for linter-detectable idioms improves performance on hard idioms. We evaluate the base model, SFT, and DPO-trained models on this benchmark.
@@
Benchmarking on Hard Idioms: We evaluate state-of-the-art open and closed-source code and reasoning LLMs on the PEP hard idiom benchmark, comparing them to MetaLint-trained models. Open-source models include instruction-tuned Qwen2.5 (Yang et al., 2024), Qwen2.5Coder (Hui et al., 2024), DeepSeek-R1-Distill-Qwen (DeepSeek-AI, 2025), Qwen3 (Team, 2025), and GPT-oss 20B/120B Agarwal et al. (2025). Closed-source models include GPT-4o (Hurst et al., 2024), o3 mini and o4 mini (OpenAI, 2025a), GPT-4.1 (OpenAI, 2025), and GPT-5 OpenAI (2025b). We select these models for their strong coding and reasoning performance and also evaluate the effects of code-specific pre-training, model scale (3B–120B), and test-time compute for open-source models.
@@ 5 Results
To test whether MetaLint training leads to cross-idiom generalization instead of mere memorization of the training idioms and whether it can produce models that can keep up with evolving code quality standards, we present the transfer performance on the synthetic data for ``easy'' idioms in section 5.1. Then we explore the extent to which MetaLint training achieves easy-to-hard generalization from the synthetic easy idioms to hard, manually curated PEP idioms in section 5.2. Finally, we compare MetaLint trained models against state-of-the-art code and reasoning models on the manually curated hard PEP idioms in section 5.3.
5.1 Generalization on Synthetic Data
Python Ruff Idioms: The performance of Qwen3-4B with and without reasoning and Llama3.2-3B-Instruct when trained on synthetic Ruff idioms and evaluated on the Ruff synthetic test set with varying transfer settings (section 4.2) is shown in Table 1 (full results in Table 18). While Table 1 shows the overall performance, we also analyze the performance broken down by each transfer setting in Table 16. The results show that the SFT stage leads to modest gains in detection and localization performance in most cases (except for a detection recall drop in the case of Llama3.2-3B-Instruct), but the DPO stage leads to huge gains in detection recall, F-score, and all localization metrics at the cost of a slight drop in detection precision. We identify that the drop in precision in the DPO stage is tightly controlled by the fraction of cases with no violations used in the DPO training and explore it in detail in Appendix D.3. Additionally, Table 16 shows that while SFT can lead to slight gains for the transfer settings (near transfer and far transfer), most gains emerge in the DPO stage, especially for non-reasoning models and detection recall. Overall this suggests that SFT can lead to memorization of the training idioms while DPO leads to generalization to novel idioms.